the sides of a rectangle are in the ratio 5:7 and the perimeter of the rectangle is 96 cm. What is the area of the rectangle?
Answer by checkley77(12844)

P=2L+2W
P=2(L+W)
96=2(5n+7n)
96=2*12n
96=24n
n=96/24
n=4 ans. for the multiplier of the ratio values.
5*4=20
7*4=28
Area=20*28=560 cm^2.